OpenStudy (anonymous):

Writing the slope-intercept? 1. Write the slope-intercept form of the line which passes through (4,2) with a slope of 1/2 a. y=1/2 x+2 b. y=1/2x c. y=1/2 x+4 2. Write the slope-intercept equation for the line through (3,7) & (9,3) a. y=-3/2x+7 b. y=-2/3x+9 c. y=3/2x+3 3. Write the slope-intercept form of the equation of the line parallel to y=5x+2, which passes through the point (4,9). a. y=5x-11 b. -5x+9 c. y=-1/5-11 Please, help me.

OpenStudy (anonymous):

Is there a formula that I'm supposed to use to figure this out?

OpenStudy (anonymous):

What numbers do I put in that? Like can you show me?

OpenStudy (anonymous):

I got y=12x+4

OpenStudy (anonymous):

oops y=1/2x+4

OpenStudy (anonymous):

Oh! So I 1/2 the 4 to get 2?

OpenStudy (anonymous):

x and y are (4,2)

OpenStudy (anonymous):

So, it's y = 1/2 x + 4

OpenStudy (anonymous):

I got y=1/2x+2

OpenStudy (anonymous):

That's one of the answer choices

OpenStudy (anonymous):

Is that correct, or...?

OpenStudy (amigatour):

it's wrong.. it's y=mx+b

OpenStudy (amigatour):

sorry, hang on.

OpenStudy (anonymous):

Okay...

OpenStudy (amigatour):

Plug the (4,2) y=mx+b to get the b.

OpenStudy (anonymous):

Okay... 2=1/2(4)+b

OpenStudy (amigatour):

sorry for confusing you/

OpenStudy (amigatour):

keep going

OpenStudy (anonymous):

It's fine, but I'm counfused because what you gave me doesn't match my answer choices

OpenStudy (amigatour):

it was for point-slope

OpenStudy (anonymous):

So... it is y=1/2x? For number 1?

OpenStudy (amigatour):

yes

OpenStudy (anonymous):

OH, okay, can you help me on the next 2?

OpenStudy (anonymous):

On number 2, I got y-1=3(x+2), is that right?

OpenStudy (anonymous):

Oops, total wrong problem

OpenStudy (anonymous):

I mean.....

OpenStudy (anonymous):

is it still y=mx+b?

OpenStudy (anonymous):

OR another formula?

OpenStudy (amigatour):

First, find the slope: m=y2-y1/x2-x1

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I got m=3-7/9-3

OpenStudy (amigatour):

Your two ordered pairs are (3,7) and (9,3). Your x1=3, x2=9 your y1=7, y2=3

OpenStudy (amigatour):

evaluate it to get your slope.

OpenStudy (anonymous):

Okay

OpenStudy (amigatour):

Don't worry, you will get used to this...you're helping me to remember all of these problems again.

OpenStudy (anonymous):

m= -4/6

OpenStudy (amigatour):

when you get your slope, use the point-slope formula: y-y1=m(x-x1)

OpenStudy (anonymous):

Okay, so I'm confused with #2 is the slope -4/6?

OpenStudy (amigatour):

Then evaluate to get your slope-intercept

OpenStudy (amigatour):

yes.

OpenStudy (amigatour):

you can choose from your two ordered pairs.

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

I chose the first on and I got 7=-4/5(3)+b

OpenStudy (amigatour):

after that, you will use y=mx+b to get your slope-interecept form of the line

OpenStudy (anonymous):

Okay, is it y=3/2 +3?

OpenStudy (amigatour):

you're doing fine.. just keep going until you have your b=?

OpenStudy (anonymous):

Oh, wait, I'm confused, how do I find b?

OpenStudy (amigatour):

olyviaevans Medals 0 I chose the first on and I got 7=-4/5(3)+b

OpenStudy (amigatour):

you finish it. remember to isolate your variable.

OpenStudy (anonymous):

Oh, okay. I got y=-2/3x+9

OpenStudy (amigatour):

Did you realize your mistake on your slope?

OpenStudy (amigatour):

m= -4/6= -2/3

OpenStudy (anonymous):

No, what mistake?

OpenStudy (anonymous):

Oh

OpenStudy (amigatour):

when you write -4/5, i was waiting for you to find it out yourself. :)

OpenStudy (anonymous):

Oh, I didn't catch it.

OpenStudy (amigatour):

just keep practicing.. it will come in handy.

OpenStudy (anonymous):

OK, thanks.

OpenStudy (amigatour):

yw